OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a vertex at the origin and a focus at (0, -4).

OpenStudy (calculusxy):

Standard form is: \(ax^2 + bx + c\) What I like to do is put this in vertex form \(a(x - h) + k\)

OpenStudy (calculusxy):

Sorry Vertex Form is \(a(x - h)^2 + k\)

OpenStudy (anonymous):

Answer choices: y = -1/4x^2 y^2 = -4x y^2 = -16x y = -1/16x^2

OpenStudy (calculusxy):

@Nnesha

OpenStudy (anonymous):

My online lessons do such a poor job at explaining most things...

OpenStudy (animelover91804):

ok... so a vertex form for the parabola is (x - h)^ = 4a(y - k) (h, k) is the vertex and a is the focal length you know the vertex is (0,0) so you have x^2 = 4ay substitute the focal length to get the equation so the answer would be.....1/36x^2=y

OpenStudy (anonymous):

You just copy and pasted that from a yahoo answers page.

OpenStudy (animelover91804):

I would so laugh at you right now... No i didn't copy it from Yahoo, Yahoo is pointless to go to bc everything is always wrong

OpenStudy (anonymous):

or another person's openstudy question because I was going through some pages and saw that word for word.

OpenStudy (anonymous):

@Nnesha Please help

Nnesha (nnesha):

alright so the equation is \[\large\rm (x-h)^2=4p(y-k)\] where (h,k) is the vertex and p is the distance between vertex and focus

Nnesha (nnesha):

 vertex is at origin so origin =(0,0)

OpenStudy (anonymous):

okay so h=0 k=0

Nnesha (nnesha):

that's correct

OpenStudy (anonymous):

Which would then mean... (x-0)^2 = 4p(y-0) x^2=4p(y)

Nnesha (nnesha):

now i'm going to draw a focus point  the focus point is below the vertex point so it should open downwards